A random sample of 225 adults was given an IQ test. It was found that 130 of them scored higher than 100. Based on this, compute a 95% confidence interval for the proportion of all adults whose IQ score is greater than 100. Then complete the table below.
Answer #1

solution: Given that, n=225 - x - 130 let ê be the sample proportion PER 130 - 0.578 2 2 Now, c = 95% = 0.95 dal-C = 1-0.95 <